2D-IR VIBRATIONAL ECHO SPECTROSCOPY OF CONDUCTING POLYMERS

\underline{AUDREY A. EIGNER}, PATRICK E. KONOLD, AARON M. MASSARI, University of Minnesota, Department of Chemistry, 207 Pleasant St SE, Minneapolis MN 55455.

2D-IR vibrational echo spectroscopy was used to examine the structural dynamics of various conducting polymers. A metal carbonyl was embedded in the polymer samples to serve as a global reporter of dynamics in these polymer thin films. Data taken on samples with metal carbonyl reporter species dissolved in NMP and embedded in polymer thin films will be presented. From these data, we are able to describe the time scales of structural motions present in these materials.